### v1

@@ Appendix A Orphan galaxies and disrupted subhaloes
Dark matter simulations produce large populations of fully disrupted subhaloes: subhaloes whose mass has fallen below the resolution of the simulation. Consequently, semi-analytic models introduce the concept of “orphan” galaxies: fully disrupted dark matter subhaloes whose stellar content remains bound (e.g., Édouard Tollet et al. 2017; Han et al. 2016). However it is now quite clear that a majority, if not all, of the disruption is numerical rather than physical in origin (van den Bosch & Ogiya 2018; van den Bosch et al. 2018). Baryons do not seem to alleviate this: Bahé et al. (2019) have shown that subhaloes are disrupted essentially just as often in the Hydrangea simulations (a set of zoom-in simulations run with the same hydrodynamics model as EAGLE), and if a dark matter subhalo is disrupted so is the galaxy. In other words, orphan galaxies essentially do not exist in the Hydrangea simulation. Hydrangea includes clusters with $M_{\mathrm{200m}}>10^{14}\,\mathrm{M}_{\odot}$; by exploring these issues in EAGLE we therefore provide an extension of those results to lower cluster mass here.
We show the fraction of disrupted satellites in our cluster sample in Figure 14, as a function of maximal subhalo mass. Here, we define a subhalo as disrupted if it has a total of 20 bound particles or less, considering all species. The disrupted fraction increases noticeably with decreasing maximal mass but even subhaloes with maximal masses $>10^{12}\,\mathrm{M}_{\odot}$ have a nonnegligible probability of being disrupted. Overall, the fraction of disrupted subhaloes is consistent with the fraction typical of dark matter simulations (Han et al. 2016), despite a lower fraction at the very high mass end which could be dominated by recent infallers.
Consistently with Bahé et al. (2019), we find little evidence for orphan galaxies. There is only one galaxy in our sample of satellites with less than 20 dark matter particles (it has 18) and only seven with less than 100 dark matter particles ($m_{\mathrm{DM}}<10^{9}\,\mathrm{M}_{\odot}$). Even at lower stellar masses there are hardly any galaxies deficient in dark matter, with only 18 galaxies (out of 28,937) with $m_{\star}>10^{7}\,\mathrm{M}_{\odot}$ and less than 20 dark matter particles ($m_{\mathrm{DM}}<2\times 10^{8}\,\mathrm{M}_{\odot}$). Only four satellite galaxy (residing in $M_{\mathrm{200m}}>10^{13}\,\mathrm{M}_{\odot}$ clusters) have lost all of their dark matter, retaining $\leq 1$ dark matter particles. The most massive of these four has a stellar mass $m_{\star}=5\times 10^{7}\,\mathrm{M}_{\odot}$ and therefore none of them are included in the analyses in this paper. These results are also consistent with Jing et al. (2019) who found only $\sim 3\%$ of satellites in EAGLE clusters have a dark matter fraction $<50\%$.
@@
Figure 14: Top panel: Unevolved subhalo mass function for present-day satellites in $M_{\mathrm{200m}}>10^{13}\,\mathrm{M}_{\odot}$ clusters. We show all subhaloes (blue) and disrupted subhaloes, defined as subhaloes which at present have less than 20 particles in total. Bottom panel: fraction of disrupted subhaloes as a function of maximal subhalo mass. The black line shows the differential fraction and the grey line shows the cumulative fraction of disrupted subhaloes above a certain mass. The dashed line marks a fraction of 0.5 for reference. All curves have been smoothed with Gaussian kernel density estimates with $\sigma=0.1$ dex in mass. Note that the lower mass limit approaches the numerical resolution of the simulation.
Appendix B Preprocessing and postprocessing of satellites in low-mass clusters

### v2

@@ Appendix A Orphan galaxies and disrupted subhaloes
Dark matter simulations produce large populations of fully disrupted subhaloes: subhaloes whose mass has fallen below the resolution of the simulation. Consequently, semi-analytic models introduce the concept of “orphan” galaxies: fully disrupted dark matter subhaloes whose stellar content remains bound (e.g. Édouard Tollet et al. 2017; Han et al. 2016). However it is now quite clear that a majority, if not all, of the disruption is numerical rather than physical in origin (van den Bosch & Ogiya 2018; van den Bosch et al. 2018). Baryons do not seem to alleviate this: Bahé et al. (2019) have shown that subhaloes are disrupted essentially just as often in the Hydrangea simulations (a set of zoom-in simulations run with the same hydrodynamics model as EAGLE), and if a dark matter subhalo is disrupted so is the galaxy. In other words, orphan galaxies essentially do not exist in the Hydrangea simulation. Hydrangea includes clusters with $M_{\mathrm{200m}}>10^{14}\,\mathrm{M}_{\odot}$; by exploring these issues in EAGLE we therefore provide an extension of those results to lower cluster mass here.
We show the fraction of disrupted satellites in our cluster sample in Figure 14, as a function of maximum subhalo mass. Here, we define a subhalo as disrupted if it has a total of 20 bound particles or fewer, considering all species. The disrupted fraction increases noticeably with decreasing maximum mass but even subhaloes with maximum masses $>10^{12}\,\mathrm{M}_{\odot}$ have a nonnegligible probability of being disrupted. Overall, the fraction of disrupted subhaloes is consistent with the fraction typical of dark matter simulations (Han et al. 2016), despite a lower fraction at the very high mass end which could be dominated by recent infallers.
Consistently with Bahé et al. (2019), we find little evidence for orphan galaxies. There is only one galaxy in our sample of satellites with fewer than 20 dark matter particles (it has 18) and only seven with less than 100 dark matter particles ($m_{\mathrm{DM}}<10^{9}\,\mathrm{M}_{\odot}$). Even at lower stellar masses there are hardly any galaxies deficient in dark matter, with only 18 galaxies (out of 28,937) with $m_{\star}>10^{7}\,\mathrm{M}_{\odot}$ and less than 20 dark matter particles ($m_{\mathrm{DM}}<2\times 10^{8}\,\mathrm{M}_{\odot}$). Only four satellite galaxy (residing in $M_{\mathrm{200m}}>10^{13}\,\mathrm{M}_{\odot}$ clusters) have lost all of their dark matter, retaining $\leq 1$ dark matter particles. The most massive of these four has a stellar mass $m_{\star}=5\times 10^{7}\,\mathrm{M}_{\odot}$ and therefore none of them are included in the analyses in this paper. These results are also consistent with Jing et al. (2019) who found only $\sim 3\%$ of satellites in EAGLE clusters have a dark matter fraction $<50\%$.
@@
Figure 14: Census of disrupted subhaloes. Top panel: Unevolved subhalo mass function for present-day satellites in $M_{\mathrm{200m}}>10^{13}\,\mathrm{M}_{\odot}$ clusters. Shown are all subhaloes (blue) and disrupted subhaloes, defined as subhaloes which at present have fewer than 20 particles in total. Bottom panel: Fraction of disrupted subhaloes as a function of maximum subhalo mass. The black line shows the differential fraction and the grey line shows the cumulative fraction of disrupted subhaloes above a certain mass. The dashed line marks a fraction of 0.5 for reference. All curves have been smoothed with Gaussian kernel density estimates with $\sigma=0.1$ dex in mass. The lower mass limit approaches the numerical resolution of the simulation.
Appendix B Pre-processing and post-processing of satellites in low-mass clusters